OpenStudy (anonymous):

i need step by step in solving for X-Y 6x+2y=22

OpenStudy (lgbasallote):

...is X-Y included in the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I mean No

OpenStudy (lgbasallote):

...well you cant solve for x and y without a second equation

OpenStudy (anonymous):

ok so it needs another equation how about 3x-10y=-12

OpenStudy (lgbasallote):

you do systems of linear equations... 6x + 2y = 22 3x - 10y = -12 multiply the second equation by -2 6x + 2y = 22 -6x +20y = 24 22y = 46 11y = 23 y = 23/11 6x +46/11 = 22 6x = 22 - 46/11 6x = 196/11 x = 98/33

jhonyy9 (jhonyy9):

6x+2y=22 so to solve this for x and y you need again one equation with x and y - than if you have this second equation on first step you see that from 6,2 and 22 you can extracting the 2 like common factor so than from this equation you will get 2(3x+y)=2(11) ,so than you can divide both sides by 2 and will remain 3x+y=11 from this you will get that y=11-3x and now if you have the second equation so than this y=11-3x you need just to substitute inside second equation in the place of y and than will get one equation just with x ,so from this you can calcule the x and after you get the x value with this you can calcule easy the term of y ... hope so much that is understandably sure and easy good luck bye